OpenStudy (anonymous):

Find value of tan when sin(56) = 0.8290 and cos(56) = 0.5592

Directrix (directrix):

Are you looking to find the tan(56)?

OpenStudy (anonymous):

it just asks for tan of theta, i tried tan of 56 and i got the question wrong

Directrix (directrix):

Assumption: The question is to find tan(56) if sin(56) = 0.8290 and cos(56) = 0.5592. Note: It would be helpful if you would post the problem as it appears in your text. You have to know your trig ratio identities. Tan(x) = sin(x) over cos(x) So, Tan(56) = Sin(56)/Cos(56) Tan (56) = .8290/.5592 Your Task: Divide .8290 by .5592 and post what you get.

OpenStudy (anonymous):

\[Use the below values \to find \tan(\theta) . Round your answer \to three decimals. \] \[\sin (56) \approx 0.8290 \] \[\cos(56) \approx 0.5592\] That's the question. I got 1.482 but i did choose that before and got it wrong

Directrix (directrix):

I got 1.482. The problem I see is that the angle theta is not defined. I assumed it is 56 degrees which is what the problem seems to insinuate. If there is a diagram of a triangle, then theta might be another angle in the triangle that is determined by using the tangent of 56 degrees. I just don't know because the problem, as posted, is poorly written. That said, we might be able to figure out the answer but it would be a guess. What are the other answer options?

OpenStudy (anonymous):

A. 1.482 B. 1.206 C. 1.788 D. 1.483 E. 0.675

Directrix (directrix):

We know A is wrong. If you crank out the tangent of 56 degrees on the calculator without looking at the given information, you get this: 1.482561 1.482 5 61 The answer is to be given to three decimals which is the nearest thousandth. Because the "5" after the 2 in 1.482 is 5 or more, the "2" in 1.482 is rounded up to a "3" That gives 1.483 as the answer to three decimal places. Yet, those other values we were given took us to 1.482 even after I checked the 4th digit. So, the question is bad. I don't know for sure because of how this problem is written but I thinkg what the answer program wants to see is --> 1.483

Directrix (directrix):

@cherub94